OpenStudy (anonymous):

Solve the equation: x2 - 12x = -36 is (6x-6)(6x+6) rite?

OpenStudy (apoorvk):

nope.! simplify what you got and see if it is equal to the original equation!

OpenStudy (ash2326):

it's close it is given \[x^2-12x=-36\] Add 36 to both the sides \[x^2-12x+36=0\] We know \[(a-b)^2=a^2-2ab+b^2\] so here a^2=x^2 b^2=36 => b=6 -2ab=-2x6=-12x so \[x^2-12x+36=(x-6)^2 \] or \[x^2-12x+36=(x-6)(x-6)\]